# Proactive Local-Minima-Free Robot Navigation: Blending Motion Prediction with Safe Control

### v1

@@ III-C Modulated Control Barrier Functions
Many safe reactive controllers, including the standard CBF-QPs, suffer from local minima or saddle points [8, 11, 18]. To generate safe inputs that ensure no local minima, we use the modulation based on-Manifold Modulated CBF-QP (MCBF-QP) [27]. Control inputs synthesized by CBF-QPs inevitably result in undesired stable equilibria even for fully-actuated robots [18, 27], which occur when $\dot{\bm{x}}=0$. A robot trapped at a undesired equilibrium will reduce the efficiency of a planned robot trajectory and even prevent it from reaching the goal.
@@ IV-B Adaptive Modulated Control
During the navigation, the virtual obstacles defined as the EFRS for pedestrians can overlap with other physical obstacles in the environment, thereby blocking certain passways to the target. When such scenarios occur, computing the exit strategy $\phi(\bm{x},h_{o})$ with respect to either the predicted reachable space or nearby physical obstacles may lead the robot to freeze at the intersecting region. Therefore, whenever a predicted reachable space of a dynamic obstacle is close enough to the robot, we instead generate $\phi(\bm{x},\bar{h})$ using the combined barrier function from all obstacles within distance $b$ to the robot. Given a set of obstacles $Q=\{o|h_{o}(\bm{x},\bm{x}_{o})\leq b,o\in O\}$, the combined single obstacle representation can be computed [8] as
@@
Constructing virtual obstacles using EFRSs introduces obstacle deformation. For instance, when a human accelerates or decelerates, their EFRS may expand or shrink drastically. MCBF-QP in [27] selects the exit tangent vector $\phi(\bm{x},\bar{h})$ using manually tuned constant parameters for iteration number $N$ and step size $\beta$. However, using constant $N$ and $\beta$ causes over-wrapping and under-wrapping issues and outputs inefficient exit strategies $\phi(\bm{x},\bar{h})$ during the shrinkage or expansion of the EFRS, as shown in Fig. 3. What’s worse, if the exit strategies selected switch frequently between opposite directions due to improper wrapping, we risk introducing local minima that MCBF-QP intends to eliminate back into the trajectories planned. Therefore, to resolve the improper wrapping issue, we propose the autonomous parameter-selection scheme (Algorithm 1) that adapts in real time as the EFRS deforms. The algorithm is motivated by the observation that proper wrapping is achieved whenever $\beta N\!\approx\!d_{\text{geo}}$, where $d_{\text{geo}}$ is the geodesic distance between the robot position $\bm{\xi}_{\text{rob}}$ and the target position $\bm{\xi}_{\text{tar}}$. This distance is defined as the length of the shortest path lying entirely on the isosurface $\{\bm{\xi}\,|\,\bar{s}(\bm{\xi})\!=\!\bar{s}(\bm{\xi}_{\text{rob}})\}$ and connects their surface projections $\bm{\xi}^{\text{pr}}_{\text{rob}}$ and $\bm{\xi}^{\text{pr}}_{\text{tar}}$. Given the correlationship between $\beta$ and $N$, fixing one while updating the other via approximated $d_{\text{geo}}$ trivially ensures proper wrapping. In our framework, $N$ is fixed since a static iteration count enables JIT compilation of the geodesic-approximation step, yielding substantial computational speedups.
@@
3 $\mathcal{B}_{\text{all}}=\text{CONTOUR}({\color[rgb]{0,0,0}\text{EDGE}(\bar{s}(\mathcal{M}),\bar{s}(\bm{\xi}_{\text{rob}})})$;
@@ V Experiment
The proposed framework is evaluated on a differential-drive Fetch robot across kinematic simulations, Gazebo simulations, and real-world experiments. Four scenarios are designed for kinematic simulations (middle and right images of Fig. 4), featuring hospital and crowd navigation tasks. The robot is required to move from the star-marked initial position to the cross-marked target, under 10 uniformly sampled initial orientations. Each algorithm’s performance is assessed in terms of safety, efficiency, and solver feasibility, as reported in Table I. The performance of the proposed methods and baselines in hospital environments is further validated in Gazebo simulations (left image of Fig. 4). In both kinematic and Gazebo hospital simulations, i.e., Scenarios 1 to 3, the EBM-based motion predictor is used; while in the crowd navigation task (Scenario 4) and the real-world experiments, the CVM-based motion predictor is used.
@@
Barrier Function Designs: In unicycle kinematics, $\bm{\xi}=[p_{x},p_{y}]^{\top}$ and $\bm{x}=[p_{x},p_{y},\theta]^{\top}$. Thus, the barrier function $h_{o}$ for shifted unicycle model can be adopted from $s_{o}$ as
$h_{o}(\bm{x},\bm{x}_{o})=h_{o}{\color[rgb]{0,0,0}(\bm{\xi},\bm{x}_{o})}={\color[rgb]{0,0,0}s_{o}}(\bm{\xi},\bm{x}_{o}).$ (22)
@@
${\color[rgb]{0,0,0}\nabla_{\bm{\xi}}h_{o}(\bm{\xi},\bm{x}_{o})^{\top}\dot{\bm{\xi}}+\nabla_{\bm{x}_{o}}h_{o}(\bm{\xi},\bm{x}_{o})^{\top}\dot{\bm{x}}_{o}\geq-\alpha({\color[rgb]{0,0,0}h_{o}}(\bm{\xi},\bm{x}_{o})).}$
For standard model, we formulate our adaptive MCBF-QP using A-CBF $h_{\text{aug}}(\bm{x},\bm{x}_{o})$ as follows, where $w\in\mathbb{R}^{+}$ [27].
${\color[rgb]{0,0,0}h_{\text{aug}}(\bm{x},\bm{x}_{o})}={\color[rgb]{0,0,0}s_{o}}(\bm{\xi},\bm{x}_{o})+w\nabla_{\bm{\xi}}{\color[rgb]{0,0,0}s_{o}}(\bm{\xi},\bm{x}_{o})^{\top}\begin{bmatrix}\cos{\theta}\\ \sin{\theta}\end{bmatrix}$ (23)
@@
Figure 5: Real-world experiments. A Fetch robot avoids three pedestrians approaching in a U-shape formation. Full video attached.

### v2

@@ III-C Modulated Control Barrier Functions
Many safe reactive controllers, including the standard CBF-QPs, suffer from local minima or saddle points [8, 11, 18]. To generate safe inputs that ensure no local minima, we use the modulation based on-Manifold Modulated CBF-QP (MCBF-QP) [27]. Control inputs synthesized by CBF-QPs inevitably result in undesired stable equilibria even for fully-actuated robots [18, 27], which occur when $\dot{\bm{x}}=0$. A robot trapped at an undesired equilibrium will reduce the efficiency of a planned trajectory and even prevent it from reaching the goal.
@@ IV-B Adaptive Modulated Control
During the navigation, the virtual obstacles defined as the EFRS for pedestrians can overlap with other physical obstacles in the environment, thereby blocking certain passageways to the target. When such scenarios occur, computing the exit strategy $\phi(\bm{x},h_{o})$ with respect to either the predicted reachable space or nearby physical obstacles may lead the robot to freeze at the intersecting region. Therefore, whenever a predicted reachable space of a dynamic obstacle is close enough to the robot, we instead generate $\phi(\bm{x},\bar{h})$ using the combined barrier function from all obstacles within distance $b$ to the robot. Given a set of obstacles $Q=\{o|h_{o}(\bm{x},\bm{x}_{o})\leq b,o\in O\}$, the combined single obstacle representation can be computed [8] as
@@
Constructing virtual obstacles using EFRSs introduces obstacle deformation. For instance, when a human accelerates or decelerates, their EFRS may expand or shrink drastically. MCBF-QP in [27] selects the exit tangent vector $\phi(\bm{x},\bar{h})$ using manually tuned constant parameters for iteration number $N$ and step size $\beta$. However, using constant $N$ and $\beta$ causes over-wrapping and under-wrapping issues and outputs inefficient exit strategies $\phi(\bm{x},\bar{h})$ during the shrinkage or expansion of the EFRS, as shown in Fig. 3. What’s worse, if the exit strategies selected switch frequently between opposite directions due to improper wrapping, we risk introducing local minima that MCBF-QP intends to eliminate back into the trajectories planned. Therefore, to resolve the improper wrapping issue, we propose the autonomous parameter-selection scheme (Algorithm 1) that adapts in real time as the EFRS deforms. The algorithm is motivated by the observation that proper wrapping is achieved whenever $\beta N\!\approx\!d_{\text{geo}}$, where $d_{\text{geo}}$ is the geodesic distance between the robot position $\bm{\xi}_{\text{rob}}$ and the target position $\bm{\xi}_{\text{tar}}$. This distance is defined as the length of the shortest path lying entirely on the isosurface $\{\bm{\xi}\,|\,\bar{s}(\bm{\xi})\!=\!\bar{s}(\bm{\xi}_{\text{rob}})\}$ and connects their surface projections $\bm{\xi}^{\text{pr}}_{\text{rob}}$ and $\bm{\xi}^{\text{pr}}_{\text{tar}}$. Given the correlation between $\beta$ and $N$, fixing one while updating the other via approximated $d_{\text{geo}}$ trivially ensures proper wrapping. In our framework, $N$ is fixed since a static iteration count enables JIT compilation of the geodesic-approximation step, yielding substantial computational speedups.
@@
3 $\mathcal{B}_{\text{all}}=\text{CONTOUR}(\text{EDGE}(\bar{s}(\mathcal{M}),\bar{s}(\bm{\xi}_{\text{rob}}))$;
@@ V Experiment
The proposed framework is evaluated on a differential-drive Fetch robot across kinematic simulations, Gazebo simulations, and real-world experiments111 Code: https://github.com/yifanxueseas/mmp_mcbf_control Video: https://youtu.be/heagLEwfcVc . Four scenarios are designed for kinematic simulations (middle and right images of Fig. 4), featuring hospital and crowd navigation tasks. The robot is required to move from the star-marked initial position to the cross-marked target, under 10 uniformly sampled initial orientations. Each algorithm’s performance is assessed in terms of safety, efficiency, and solver feasibility, as reported in Table I. The performance of the proposed methods and baselines in hospital environments is further validated in Gazebo simulations (left image of Fig. 4). In both kinematic and Gazebo hospital simulations, i.e., Scenarios 1 to 3, the EBM-based motion predictor is used; while in the crowd navigation task (Scenario 4) and the real-world experiments, the CVM-based motion predictor is used.
@@
Barrier Function Designs: In unicycle kinematics, $\bm{\xi}=[p_{x},p_{y}]^{\top}$ and $\bm{x}=[p_{x},p_{y},\theta]^{\top}$. Thus, the barrier function $h_{o}$ for the shifted unicycle model can be adopted from $s_{o}$ as
$h_{o}(\bm{x},\bm{x}_{o})=h_{o}(\bm{\xi},\bm{x}_{o})=s_{o}(\bm{\xi},\bm{x}_{o}).$ (22)
@@
$\nabla_{\bm{\xi}}h_{o}(\bm{\xi},\bm{x}_{o})^{\top}\dot{\bm{\xi}}+\nabla_{\bm{x}_{o}}h_{o}(\bm{\xi},\bm{x}_{o})^{\top}\dot{\bm{x}}_{o}\geq-\alpha(h_{o}(\bm{\xi},\bm{x}_{o})).$
For the standard model, we formulate our adaptive MCBF-QP using A-CBF $h_{\text{aug}}(\bm{x},\bm{x}_{o})$ as follows, where $w\in\mathbb{R}^{+}$ [27].
$h_{\text{aug}}(\bm{x},\bm{x}_{o})=s_{o}(\bm{\xi},\bm{x}_{o})+w\nabla_{\bm{\xi}}s_{o}(\bm{\xi},\bm{x}_{o})^{\top}\begin{bmatrix}\cos{\theta}\\ \sin{\theta}\end{bmatrix}$ (23)
@@
Figure 5: Real-world experiments. A Fetch robot avoids three pedestrians approaching in a U-shape formation.